Use taylors formula to find quadratic and cubic approximations of f(x, y) =exsin y at origin. Estimate the error of the approximations if modx less than equal to 0.1 and mod y less than equal to 0.1

1
Expert's answer
2021-12-14T08:09:07-0500

Solution:


  • The given function is "f(x,y)=e^xsin(y)"
  • The objective is to find the quadratic approximation of a given function using the Taylor formula and also estimate the error in the approximation if "|\nx\n|\n\u2264\n0.1" and "|\ny\n|\n\u2264\n0.1\n ."


Calculate First partial derivatives at "(0,0)" .


"f\nx\n=e^xsin(y)"

"f\nx\n(\n0\n,\n0\n)\n=\n0"

"f\ny\n=\ne^xcos(y)"

"f\ny\n(\n0\n,\n0\n)\n=\n1"


Calculate second derivatives at "(0,0)"


"f\nx\nx\n=\ne^xsin(y)"

"f\nx\n(\n0\n,\n0\n)\n=\n0"

"f\ny\n=\ne^xcos(y)"

"f\ny\n(\n0\n,\n0\n)\n=\n1"

"f\nx\ny\n=\n-e^xsin(y)"

"f\nx\ny\n(\n0\n,\n0\n)\n=\n0"


Therefore quadratic approximation of "f(x,y)" at origin using Taylor formula is:


"T
(
x
,
y
)
≈
9
+
\frac{1}{2}
[
−
9
x^2
−
9
y^2
]
≈
9
−
\frac{9}{2}
[
x^2
+
y^2
]"


Next Calculating error at the origin using "|\nx\n|\n\u2264\n0.1" and "|\ny\n|\n\u2264\n0.1" as:


"E
(
x
,
y
)
≤
\frac{1}{2}
×
9
[
(
0.1
)^2
+
(
0.1
)
^2
]
≤
0.08685"


Thus, the Quadratic polynomial using the Taylor formula is "9
−
\frac{9}{2}
[
x
^2
+
y
^2
]" and error in the approximation is less than 0.08685.